OpenStudy (setsuna-yuregeshi):

The specification for a length x is 43.6 cm with a tolerance of 0.1 cm. Write the specification as an absolute value inequality.

OpenStudy (setsuna-yuregeshi):

I really need help, but on how to do it. Not just the answer please...

Vocaloid (vocaloid):

hint: the absolute value of the difference between x and 43.6 is less than or equal to the tolerance

OpenStudy (setsuna-yuregeshi):

Um... So |x-43.6|<or= 43.7? @Vocaloid

Vocaloid (vocaloid):

almost. what is the tolerance?

OpenStudy (setsuna-yuregeshi):

What does tolerance mean?

OpenStudy (setsuna-yuregeshi):

lol

OpenStudy (setsuna-yuregeshi):

I don't know :(

Vocaloid (vocaloid):

"tolerance of 0.1" means that the value of x will be no more than 43.6 + 0.1, and no less than 43.6 - 0.1 basically, the tolerance gives you an upper bound and a lower bound for the possible x values

Vocaloid (vocaloid):

so our inequality is |x - 43.6| <= 0.1

OpenStudy (setsuna-yuregeshi):

eh?

OpenStudy (setsuna-yuregeshi):

How come?

Vocaloid (vocaloid):

think about what this statement means |x - 43.6| <= 0.1 |x - 43.6| represents the distance between x and 43.6. the inequality means that x will be no more than 0.1 units away from 43.6

Vocaloid (vocaloid):



OpenStudy (setsuna-yuregeshi):

Ooh,

Vocaloid (vocaloid):